An electric field (E=4.5N/C) passes through a flat surface area (0.30m) with an angle of 45°. What will be the magnitude of the electric flux?


1
Expert's answer
2022-03-08T08:56:49-0500

The electric flux

"\Phi=EA\cos\theta\\
=\rm 4.5\:N/C*0.3\:m^2*\cos45^\circ=0.95\:N\cdot m^2/C"